OpenStudy (anonymous):

Find the derivative by the limit process: h(s)=3+(2/3)s

OpenStudy (anonymous):

h'(s) = 2/3

OpenStudy (anonymous):

f(s+a) -f(x) = 3+(2/3)(s+a) - 3-(2/3)s = (2/3)a ====> h'(s) = lim (a--->0 ) [(2/3)a/a] =2/3